The graph of the exponential function: f(x)=a^x , contains the point (3, (1/64)) Find a.
Answer by Alan3354(69443)

The graph of the exponential function: f(x)=a^x , contains the point (3, (1/64)) Find a.
------------------
1/64 = a^3 (that's what it says)
a = 1/4
-----------
If you want to graph it, get the FREE software at
www.padowan.dk.com/graph/